Solve each of the following equations for y in terms of x.
3/4(2x-4y)=9

Answer by Alan3354(21605)

3/4(2x-4y)=9
3(2x-4y) = 36
2x-4y = 12
-4y = 12-2x
y = (x/2) - 3 or (x-6)/2